OpenStudy (zubhanwc3):

help solving a limit where x approaches 0 Need help solving algebraically

OpenStudy (zubhanwc3):

The equation is \[\lim_{x \rightarrow 0} \frac{ \frac{ 1 }{ 2+x }-\frac{ 1 }{ 2 } }{ x }\]

OpenStudy (zubhanwc3):

how would i solve this algebraically?

OpenStudy (johnweldon1993):

Well focus on the numerator first \[\large \frac{1}{2 + x} - \frac{1}{2}\] What is the common denominator here?

OpenStudy (zubhanwc3):

2

OpenStudy (johnweldon1993):

Well actually is would be 2(x + 2) If we multiply the first fraction by 2...and the second fraction by (2 + x) we will have \[\large \frac{2}{2(2 + x)} - \frac{2 + x}{2(2 + x)}\] Now we can put them over the common denominator \[\large \frac{2 - (2 + x)}{2(2 + x)}\] simplify that a bit more...

OpenStudy (zubhanwc3):

2-1 over 2? so 1/2?

OpenStudy (johnweldon1993):

\[\large \frac{2 - (2 + x)}{2(x + 2)}\] distribute out the '-' sign \[\large \frac{2 - 2 - x}{2(x + 2)}\] so \[\large \frac{-x}{2(x + 2)}\] make sense there? well that was just the numerator of the original question So lets add in the 'x' we ignored before \[\Large \frac{\frac{-x}{2(x + 2)}}{x}\] what do we do next?

OpenStudy (zubhanwc3):

so u multiply by 1/x to the whole thing making it \[\frac{ -x }{ 2x(2+x) }\]

OpenStudy (zubhanwc3):

then the x cancels out, making it \[\frac{ -1 }{ 2(2+x) }\]

OpenStudy (zubhanwc3):

and by substituting, u get -1/4 correct?

OpenStudy (zubhanwc3):

@johnweldon1993 is that correct?

OpenStudy (johnweldon1993):

Correct indeed